# JAFAR: Jack up Any Feature at Any Resolution

### v1

@@ Abstract
Foundation Vision Encoders have become essential for a wide range of dense vision tasks. However, their low-resolution spatial feature outputs necessitate feature upsampling to produce the high-resolution modalities required for downstream tasks. In this work, we introduce JAFAR—a lightweight and flexible feature upsampler that enhances the spatial resolution of visual features from any Foundation Vision Encoder to an arbitrary target resolution. JAFAR employs an attention-based module designed to promote semantic alignment between high-resolution queries—derived from low-level image features—and semantically enriched low-resolution keys, using Spatial Feature Transform (SFT) modulation. Notably, despite the absence of high-resolution supervision, we demonstrate that learning at low upsampling ratios and resolutions generalizes remarkably well to significantly higher output scales. Extensive experiments show that JAFAR effectively recovers fine-grained spatial details and consistently outperforms existing feature upsampling methods across a diverse set of downstream tasks.
@@
Figure 1: JAFAR upsamples features from any foundation vision encoder to any image resolution, using the input image as high-resolution guidance. It generates sharp, boundary-aligned feature maps and serves as a versatile drop-in module for a variety of downstream tasks—including semantic segmentation, open-vocabulary segmentation, depth estimation, CAM evaluation, and bird’s-eye-view segmentation—consistently enhancing performance.
@@ 1 Introduction
Foundation vision encoders—whether trained with language supervision [1, 2, 3, 4, 5] or purely on visual data [6, 7, 8]—have become core components of modern computer vision pipelines. Vision-language models excel at tasks requiring generalization, such as zero-shot classification and open-vocabulary segmentation [9, 10]. In contrast, image-only models, which focus on visual structure, often outperform in dense prediction tasks that demand fine-grained spatial reasoning, including semantic segmentation, depth estimation, object discovery, and point tracking [11, 12, 13].
To handle high-resolution inputs and large-scale training, foundation vision encoders typically downsample spatial information aggressively—by a factor of 14×14$\times$14 × to 16×16$\times$16 ×—yielding semantically rich but spatially coarse feature maps. This compression introduces a bottleneck for downstream tasks that require pixel-level accuracy. As a result, downstream pipelines [14, 15, 11, 16, 17] often rely on interpolation or dedicated modules [18, 19] designed to produce high-resolution outputs.
Several strategies have been explored to overcome this bottleneck, but each comes with trade-offs in efficiency and output quality. A straightforward solution is to apply training-free interpolation methods, such as bilinear upsampling. While computationally efficient, these direct interpolations—relying solely on low-resolution feature maps—fail to leverage information from the original high-resolution image, often resulting in blurry outputs. Alternatively, one can upsample the input image prior to encoding to increase feature resolution. However, this approach significantly increases computational cost due to the quadratic complexity of self-attention—common in foundation models—and may introduce artifacts in the feature maps, ultimately degrading performance [20, 21].
Focusing specifically on a target downstream task, [22, 23, 24, 25, 26] learn feature upsamplers using high-resolution supervision from task-specific labels. While generally lightweight, these upsamplers depend on labeled data tied to the end application, which limits their generalization and may bias the learned features toward optimizing task-specific losses. To address this, recent methods such as LiFT [27] and FeatUp [28] adopt task-agnostic training objectives. LiFT is trained to perform 2×2$\times$2 × upsampling by regressing feature maps extracted from images at twice the input resolution. However, its convolution-based architecture is limited to fixed 2×2$\times$2 × scaling, restricting its flexibility for arbitrary output resolutions. FeatUp, in contrast, uses augmented views and self-reconstruction to support higher upsampling ratios. Yet, its Joint Bilateral Upsampling (JBU) variant suffers from over-smoothed outputs, while its implicit variant requires training the upsampler for each image, making it impractical in real-world scenarios.
@@
To enable upsampling to arbitrary target resolutions, we formulate our approach as a global interpolation mechanism using a cross-attention block. The success of this attention-based method depends critically on achieving strong semantic alignment between the queries and keys. In JAFAR, we construct these representations asymmetrically (see Fig. 2): the queries retain high-resolution, low-level details such as color and texture, while the keys are hybrid features that combine high-level semantics with spatial cues. We find that enriching the keys with low-level information significantly improves query-key alignment and enhances generalization to unseen output resolutions.
Additionally, we propose a simple training objective similar to [27], but without being constrained to a fixed upsampling factor. Notably, we find that training on low upsampling factors at low resolutions (e.g., $8\times 8\rightarrow 32\times 32$) is sufficient to generalize effectively to much larger scales (e.g., $32\times 32\rightarrow 448\times 448$) while keeping memory requirements low during training, unlike training directly at higher resolutions and factors. Our contributions can be summarized as follows:
@@ 2 Related Work
Feature upsampling aims to increase the spatial resolution of intermediate feature maps within deep networks—analogous to image upsampling, but performed in a latent space. This process is essential for dense prediction tasks such as segmentation and depth estimation, where fine spatial detail is critical. Traditional interpolation techniques, such as bilinear, spline, or Lanczos [29, 30, 31, 32], provide simple and efficient baselines but do not adapt to the underlying content. Recent neural methods improve on static approaches by learning to reconstruct high-resolution features from data. These methods fall into two categories: task-dependent, trained with downstream labels supervision, and task-agnostic, trained independently of the end task. For example, CARAFE [22] and DySample [24] predict content-aware kernels or dynamic sampling positions. SAPA [23] and ReSFU [25] exploit a similarity based approach to refine spatial semantics. However, task-specific reliance on labels limits generalization. Recent task-agnostic methods like LiFT [27] and FeatUp [28] remove this dependency. LiFT introduces a CNN module trained with a simple fixed scale training objective, while FeatUp relies on a complex multi-loss objective which makes training difficult to tune in practice. Moreover, it requires training both an upsampler and a downsampler, adding unnecessary computational overhead. Notably, its best performance is achieved through per-image optimization, further limiting its practicality. In contrast, JAFAR provides a scalable task-agnostic framework that generalizes across resolutions without complex pipelines or per-image optimization, showing strong performance even when trained on small upsampling factors at low resolution.
@@
Upsampling modules architectures vary from fixed-scale decoders to continuous resolution predictors. LiFT [27] relies on a lightweight CNN module trained to upsample by a fixed factor, making further scaling dependent on iterative use which leads to performance degradation or additional interpolation steps. FeatUp [28] introduces two architectural variants: a fast Joint Bilateral Upsampler (JBU) and a more accurate implicit network allowing continuous querying. While the implicit model yields superior results, it suffers from significant inference latency due to per-image optimization. JBU, on the other hand, trades expressivity for scalability, stacking multiple ×$\times 2$ stages to achieve higher upsampling ratios. Attention-based designs, as in SAPA [23] and ReSFU [25], offer increased flexibility by modeling affinities between features across scales. These methods exploit spatial similarities to reconstruct high-resolution maps. JAFAR innovates by unifying low- and high-resolution streams: it aligns high-resolution queries and low-resolution keys using shared low-level features while enriching the representation with additional semantic cues. This design maintains spatial alignment and expressivity even at large upsampling ratios, offering a robust and scalable architecture for feature reconstruction.
@@
Feature modulation techniques modulate features using conditioning information, thereby enabling spatially or semantically guided transformations. Early forms such as Conditional BatchNorm [33], AdaIN [34], and FiLM [35] apply learned scale ($\gamma$) and shift ($\beta$) parameters per channel, derived from global conditioning signals. These methods are effective for tasks involving global transformations like style transfer or classification. However, their spatial invariance limits expressiveness in tasks requiring spatial sensitivity. SPADE [36] and SFT [37] address this limitation by computing $\gamma$ and $\beta$ as full-resolution maps conditioned on dense inputs like segmentation masks. This spatial adaptation improves expressivity by allowing each feature location to be modulated uniquely. Moreover, this form of modulation can be interpreted as a parameterized, learned recombination of feature channels, similar to $1\times 1$ convolutions but more powerful due to the spatial specificity. In JAFAR, modulation is employed not merely to shift feature distributions but to enable semantically enriched reconstruction by injecting high-resolution semantics directly into the upsampling pipeline. This allows a richer linear combination of features, improving generalization and spatial expressivity without relying on per-pixel optimization at test time [28].
@@ 3.1 Architecture
The overall flow of our architecture is illustrated in Fig. 2. JAFAR takes as input a high-resolution image $I\in\mathbb{R}^{3\times H\times W}$ and a low-resolution feature map $F_{lr}=f(I)\in\mathbb{R}^{C\times h_{k}\times w_{k}}$, extracted from a frozen vision encoder $f$. The image $I$ is first projected into a higher-dimensional space and processed by a lightweight encoder $E_{\theta}$ to obtain an intermediate representation $I_{E}=E_{\theta}(I)\in\mathbb{R}^{d\times H\times W}$, further enriched with RoPE positional embeddings [38].
@@
Relying exclusively on low-resolution features from the vision encoder to construct keys leads to poor generalization and noticeable artifacts, primarily due to an abstraction gap between these coarse features and the fine-grained queries. As demonstrated in Sec. 4, this mismatch results in inconsistent alignment across resolutions. To address this issue, we construct hybrid key representations that retain structural alignment with the queries while incorporating the semantic richness of the vision encoder. This is achieved by encoding the intermediate representation $I_{E}$ to produce $I_{K}$, which is then downsampled to match the spatial resolution of the encoder’s feature map to produce preliminary keys $\tilde{K}$. These are further modulated using the encoder feature map $F_{lr}\in\mathbb{R}^{C\times h_{k}\times w_{k}}$ through a spatial semantic feature modulation inspired by [36, 37]:
$K=\gamma_{F}\cdot\tilde{K}+\beta_{F},$ (2)
@@
To perform upsampling, we use a simplified attention mechanism where attention weights are computed via a scaled dot product between queries and semantically modulated keys. Crucially, both queries and keys have been been enriched with relative positional embeddings using RoPE  [38], which introduces an inductive bias that captures spatial relationships between queries and keys. This positional encoding allows us to entirely bypass the arbitrary selection of neighboring keys for each query, a common heuristic in prior similarity-based methods such as [23, 25]. Without this positional grounding, the attention mechanism lacks spatial awareness and generalizes poorly to unseen resolutions. In practice, we use multiple attention heads to increase expressivity and average the resulting attention weights across heads after applying softmax. The resulting attention map $A$ is then used to interpolate the low-resolution encoder features $F_{lr}$ via a simple matrix product: $\hat{F}_{HR}=A\cdot F_{lr}$. By avoiding a learned value projection, we preserve the original feature content and enable a resolution-agnostic design that generalizes reliably across scales.
@@ 3.2 Training Pipeline
To enable this, we introduce a fully annotation-free training scheme that relies only on multi-resolution views of the same image, easily obtained through standard downsampling. Given a high-resolution image $I_{HR}\in\mathbb{R}^{3\times H\times W}$, we generate a downsampled version $I_{LR}\in\mathbb{R}^{3\times\lfloor\frac{H}{\delta}\rfloor\times\lfloor\frac{W}{\delta}\rfloor}$ using a randomly sampled factor $\delta\in[2,4]$. Both images are passed through the frozen vision encoder $f$, producing two feature maps: $F_{hr}=f(I_{HR})\in\mathbb{R}^{C\times h\times w}$ and $F_{lr}=f(I_{LR})\in\mathbb{R}^{C\times\lfloor\frac{h}{\delta}\rfloor\times\lfloor\frac{w}{\delta}\rfloor}$, respectively. JAFAR then takes $I_{HR}$ and $F_{lr}$ as input to predict an upsampled feature map $\hat{F}_{hr}$. The predicted output is aligned with the target $F_{hr}$ using a simple alignment loss, which combines cosine similarity and L2 distance [20]:
$\mathcal{L}(\hat{F}_{hr},F_{hr})=1-\text{cos}(\hat{F}_{hr},F_{hr})+||\hat{F}_{hr}-F_{hr}||_{2}.$ (3)
Notably, during training, JAFAR is only exposed to moderate upsampling factors (up to 4×4$\times$4 ×), yet it generalizes remarkably well to much higher resolutions at test time—without access to any ground-truth high-resolution features.
@@
While our training objective is similar to that of LiFT, our approach demonstrates significantly greater capability, as shown in Tabs. 1 and 2. LiFT relies on a CNN-based architecture and is trained for fixed 2×2$\times$2 × upsampling at two predefined resolutions. As a result, it struggles to extrapolate beyond that setting without additional heuristics such as iterative upsampling or bilinear fallback. In contrast, JAFAR maintains a resolution-agnostic design which generalizes to much higher upsampling factors using this similar simple training setup.
@@ 4.1 Experimental Setup
In our experiments, we train JAFAR on a single NVIDIA A100 on ImageNet training set for 100K steps using AdamW optimizer [39], with a learning rate of $2\text{e}{-4}$ and a batch size of 4. The input images fed into the foundation vision encoder are resized to $448\times 448$, producing high-resolution target feature maps $F_{hr}$ of size $32\times 32$ or $28\times 28$, depending on the encoder’s patch size (14 or 16). For improved training efficiency, the guidance image input to JAFAR is downsampled to $224\times 224$.

### v2

@@ Abstract
Foundation Vision Encoders have become essential for a wide range of dense vision tasks. However, their low-resolution spatial feature outputs necessitate feature upsampling to produce the high-resolution modalities required for downstream tasks. In this work, we introduce JAFAR, a lightweight and flexible feature upsampler that enhances the spatial resolution of visual features from any Foundation Vision Encoder to an arbitrary target resolution. JAFAR employs an attention-based module designed to promote semantic alignment between high-resolution queries, derived from low-level image features, and semantically enriched low-resolution keys, using Spatial Feature Transform (SFT) modulation. Notably, despite the absence of high-resolution supervision, we demonstrate that learning at low upsampling ratios and resolutions generalizes remarkably well to significantly higher output scales. Extensive experiments show that JAFAR effectively recovers fine-grained spatial details and consistently outperforms existing feature upsampling methods across a diverse set of downstream tasks.
@@
Figure 1: JAFAR upsamples features from any foundation vision encoder to any image resolution, using the input image as high-resolution guidance. It generates sharp, boundary-aligned feature maps and serves as a versatile drop-in module for a variety of downstream tasks, including semantic segmentation, open-vocabulary segmentation, depth estimation, CAM evaluation, and bird’s-eye-view segmentation—consistently enhancing performance.
@@ 1 Introduction
Foundation vision encoders—whether trained with language supervision [clip, siglip, siglip2, tulip, radiov2.5] or purely on visual data [dino, dinov2, capi]—have become core components of modern computer vision pipelines. Vision-language models excel at tasks requiring generalization, such as zero-shot classification and open-vocabulary segmentation [maskclip, sclip]. In contrast, image-only models, which focus on visual structure, often outperform in dense prediction tasks that demand fine-grained spatial reasoning, including semantic segmentation, depth estimation, object discovery, and point tracking [depth_anything, dino_meets_text, dinotracker].
To handle high-resolution inputs and large-scale training, foundation vision encoders typically downsample spatial information aggressively—by a factor of $14\times$ to $16\times$—yielding semantically rich but spatially coarse feature maps. This compression introduces a bottleneck for downstream tasks that require pixel-level accuracy. As a result, downstream pipelines [detr, odise, depth_anything, depth_anythingv2, diffcut] often rely on interpolation or dedicated modules [dpt, mask2former] designed to produce high-resolution outputs.
Several strategies have been explored to overcome this bottleneck, but each comes with trade-offs in efficiency and output quality. A straightforward solution is to apply training-free interpolation methods, such as bilinear upsampling. While computationally efficient, these direct interpolations—relying solely on low-resolution feature maps—fail to leverage information from the original high-resolution image, often resulting in blurry outputs. Alternatively, one can upsample the input image prior to encoding to increase feature resolution. However, this approach significantly increases computational cost due to the quadratic complexity of self-attention—common in foundation models—and may introduce artifacts in the feature maps, ultimately degrading performance [dvt, vitar].
Focusing specifically on a target downstream task, [carafe, sapa, dysample, resfu, FADE] learn feature upsamplers using high-resolution supervision from task-specific labels. While generally lightweight, these upsamplers depend on labeled data tied to the end application, which limits their generalization and may bias the learned features toward optimizing task-specific losses. To address this, recent methods such as LiFT [lift] and FeatUp [featup] adopt task-agnostic training objectives. LiFT is trained to perform $2\times$ upsampling by regressing feature maps extracted from images at twice the input resolution. However, its convolution-based architecture is limited to fixed $2\times$ scaling, restricting its flexibility for arbitrary output resolutions. FeatUp, in contrast, uses augmented views and self-reconstruction to support higher upsampling ratios. Yet, its Joint Bilateral Upsampling (JBU) variant suffers from over-smoothed outputs, while its implicit variant requires training the upsampler for each image, making it impractical in real-world scenarios.
@@
To enable upsampling to arbitrary target resolutions, we formulate our approach as a global interpolation mechanism using a cross-attention block. The success of this attention-based method depends critically on achieving strong semantic alignment between the queries and keys. In JAFAR, we construct these representations asymmetrically (see Fig.˜2): the queries retain high-resolution, low-level details such as color and texture, while the keys are hybrid features that combine high-level semantics with spatial cues. We find that enriching the keys with low-level information significantly improves query-key alignment and enhances generalization to unseen output resolutions.
Additionally, we propose a simple training objective similar to [lift], but without being constrained to a fixed upsampling factor. Notably, we find that training on low upsampling factors at low resolutions (e.g., $8\times 8\rightarrow 32\times 32$) is sufficient to generalize effectively to much larger scales (e.g., $32\times 32\rightarrow 448\times 448$) while keeping memory requirements low during training, unlike training directly at higher resolutions and factors. Our contributions can be summarized as follows:
@@ 2 Related Work
Feature upsampling aims to increase the spatial resolution of intermediate feature maps within deep networks—analogous to image upsampling, but performed in a latent space. This process is essential for dense prediction tasks such as segmentation and depth estimation, where fine spatial detail is critical. Traditional interpolation techniques, such as bilinear, spline, or Lanczos [maeland1988comparison, lanczos, cardinalspline, cubic], provide simple and efficient baselines but do not adapt to the underlying content. Recent neural methods improve on static approaches by learning to reconstruct high-resolution features from data. These methods fall into two categories: task-dependent, trained with downstream labels supervision, and task-agnostic, trained independently of the end task. For example, CARAFE [carafe] and DySample [dysample] predict content-aware kernels or dynamic sampling positions. SAPA [sapa] and ReSFU [resfu] exploit a similarity based approach to refine spatial semantics. However, task-specific reliance on labels limits generalization. Recent task-agnostic methods like LiFT [lift] and FeatUp [featup] remove this dependency. LiFT introduces a CNN module trained with a simple fixed scale training objective, while FeatUp relies on a complex multi-loss objective which makes training difficult to tune in practice. Moreover, it requires training both an upsampler and a downsampler, adding unnecessary computational overhead. Notably, its best performance is achieved through per-image optimization, further limiting its practicality. In contrast, JAFAR provides a scalable task-agnostic framework that generalizes across resolutions without complex pipelines or per-image optimization, showing strong performance even when trained on small upsampling factors at low resolution.
@@
Upsampling modules architectures vary from fixed-scale decoders to continuous resolution predictors. LiFT [lift] relies on a lightweight CNN module trained to upsample by a fixed factor, making further scaling dependent on iterative use which leads to performance degradation or additional interpolation steps. FeatUp [featup] introduces two architectural variants: a fast Joint Bilateral Upsampler (JBU) and a more accurate implicit network allowing continuous querying. While the implicit model yields superior results, it suffers from significant inference latency due to per-image optimization. JBU, on the other hand, trades expressivity for scalability, stacking multiple $\times 2$ stages to achieve higher upsampling ratios. Attention-based designs, as in SAPA [sapa] and ReSFU [resfu], offer increased flexibility by modeling affinities between features across scales. These methods exploit spatial similarities to reconstruct high-resolution maps. JAFAR innovates by unifying low- and high-resolution streams: it aligns high-resolution queries and low-resolution keys using shared low-level features while enriching the representation with additional semantic cues. This design maintains spatial alignment and expressivity even at large upsampling ratios, offering a robust and scalable architecture for feature reconstruction.
@@
Feature modulation techniques modulate features using conditioning information, thereby enabling spatially or semantically guided transformations. Early forms such as Conditional BatchNorm [condbatchnorm], AdaIN [adain], and FiLM [film] apply learned scale ($\gamma$) and shift ($\beta$) parameters per channel, derived from global conditioning signals. These methods are effective for tasks involving global transformations like style transfer or classification. However, their spatial invariance limits expressiveness in tasks requiring spatial sensitivity. SPADE [spade] and SFT [sft] address this limitation by computing $\gamma$ and $\beta$ as full-resolution maps conditioned on dense inputs like segmentation masks. This spatial modulation enhances expressivity by enabling unique adjustments at each feature location. It can be interpreted as a parameterized, learned recombination of feature channels, analogous to a $1\times 1$ convolution but extended with spatially varying weights. In JAFAR, feature modulation is used not only to shift feature distributions but also to inject semantics directly into the upsampling pipeline. This provides richer linear combinations of features, improving generalization and spatial fidelity without the need for per-image optimization [featup].
@@ 3.1 Architecture
The overall flow of our architecture is illustrated in Fig.˜2. JAFAR takes as input a high-resolution image $I\in\mathbb{R}^{3\times H\times W}$ and a low-resolution feature map $F_{lr}=f(I)\in\mathbb{R}^{C\times h_{k}\times w_{k}}$, extracted from a frozen vision encoder $f$. The image $I$ is first projected into a higher-dimensional space and processed by a lightweight encoder $E_{\theta}$ to obtain an intermediate representation $I_{E}=E_{\theta}(I)\in\mathbb{R}^{d\times H\times W}$, further enriched with RoPE positional embeddings [rope].
@@
Relying exclusively on low-resolution features from the vision encoder to construct keys leads to poor generalization and noticeable artifacts, primarily due to an abstraction gap between these coarse features and the fine-grained queries. As demonstrated in Sec.˜4, this mismatch results in inconsistent alignment across resolutions. To address this issue, we construct hybrid key representations that retain structural alignment with the queries while incorporating the semantic richness of the vision encoder. This is achieved by encoding the intermediate representation $I_{E}$ to produce $I_{K}$, which is then downsampled to match the spatial resolution of the encoder’s feature map to produce preliminary keys $\tilde{K}$. These are further modulated using the vision encoder feature map $F_{lr}\in\mathbb{R}^{C\times h_{k}\times w_{k}}$ through a spatial semantic feature modulation inspired by [spade, sft]:
$K=\gamma_{F}\odot\tilde{K}+\beta_{F},$ (2)
@@
To perform upsampling, we use a simplified attention mechanism where attention weights are computed via a scaled dot product between queries and semantically modulated keys. Crucially, both queries and keys have been enriched with relative positional embeddings using RoPE [rope], which introduces an inductive bias that captures spatial relationships between queries and keys. This positional encoding allows us to entirely bypass the arbitrary selection of neighboring keys for each query, a common heuristic in prior similarity-based methods such as [sapa, resfu]. Without this positional grounding, the attention mechanism lacks spatial awareness and generalizes poorly to unseen resolutions. In practice, we use multiple attention heads to increase expressivity and average the resulting attention weights across heads after applying softmax. The resulting attention map $A$ is then used to interpolate the low-resolution encoder features $F_{lr}$ via a simple matrix product: $\hat{F}_{HR}=A\cdot F_{lr}$. By avoiding a learned value projection, we preserve the original feature content and enable a resolution-agnostic design that generalizes reliably across scales.
@@ 3.2 Training Pipeline
To enable this, we introduce a fully annotation-free training scheme that relies only on multi-resolution views of the same image, easily obtained through standard downsampling. Given a high-resolution image $I_{HR}\in\mathbb{R}^{3\times H\times W}$, we generate a downsampled version $I_{LR}\in\mathbb{R}^{3\times\lfloor\frac{H}{\delta}\rfloor\times\lfloor\frac{W}{\delta}\rfloor}$ using a randomly sampled factor $\delta\in[2,4]$. Both images are passed through the frozen vision encoder $f$, producing two feature maps: $F_{hr}=f(I_{HR})\in\mathbb{R}^{C\times h\times w}$ and $F_{lr}=f(I_{LR})\in\mathbb{R}^{C\times\lfloor\frac{h}{\delta}\rfloor\times\lfloor\frac{w}{\delta}\rfloor}$, respectively. JAFAR then takes $I_{HR}$ and $F_{lr}$ as input to predict an upsampled feature map $\hat{F}_{hr}$. The predicted output is aligned with the target $F_{hr}$ using a simple alignment loss, which combines cosine similarity and L2 distance [dvt]:
[…]
Notably, during training, JAFAR is only exposed to moderate upsampling factors (up to $4\times$), yet it generalizes remarkably well to much higher resolutions at test time—without access to any ground-truth high-resolution features.
@@
While our training objective is similar to that of LiFT, our approach demonstrates significantly greater capability, as shown in Tabs.˜1 and 2. LiFT relies on a CNN-based architecture and is trained for fixed $2\times$ upsampling at two predefined resolutions. As a result, it struggles to extrapolate beyond that setting without additional heuristics such as iterative upsampling or bilinear fallback. In contrast, JAFAR maintains a resolution-agnostic design which generalizes to much higher upsampling factors using this similar simple training setup.
@@ 4.1 Experimental Setup
In our experiments, we train JAFAR on a single NVIDIA A100 on ImageNet training set for 100K steps using AdamW optimizer [adamw], with a learning rate of $2\text{e}{-4}$ and a batch size of 4. The input images fed into the foundation vision encoder are resized to $448\times 448$, producing high-resolution target feature maps $F_{hr}$ of size $32\times 32$ or $28\times 28$, depending on the encoder’s patch size (14 or 16). For improved training efficiency, the guidance image input to JAFAR is downsampled to $224\times 224$.